b) A homogenous liquid phase reaction with the given stoichiometry and kinetics A→S,-rA=kCA^2, takes place with 50% conversion in a mixed flow reactor. Find the conversion if this reactor is replaced by another mixed flow reactor having volume 6 times that of the original reactor-all else remain unchanged. Find the conversion if the original reactor is replaced by a plug flow of the same size- all else remains unchanged.


1
Expert's answer
2021-09-10T00:04:00-0400

According to the rate of the reaction, the reaction is of second order. The conversion in a mixed flow reactor is, "X_a=0.5" ForFor a mixed flow reactor, the general reactor design equation is "\frac{V}{F_{A0}}=\frac{X_a}{r_a}"

where V is the volume of the reactor, "F_{A0}"  is volumetric flow rate, "C_{A0}"  is the initial concentration of A, and  −r is the rate of the equation. In Plug flow reactor,"\frac{V}{F_{A0}}=\frac{dX_a}{r_a}" where all the terms have similar notations to mixed flow reactor. 




a) The volume of the new reactor is 6 times the volume of the original reactor, V'=6×V

"V'=6\u00d7V"

Therefore,





The conversion can never be greater than 1. So, the conversion is 0.75.


b) In the case of PFR,"\frac{V}{F_{A0}}=\frac{dX_a}{r_a}"



a) Conversion is 75%b) Conversion is 66.66%